OpenStudy (anonymous):

for the given function determine consecutive values of x between which each real zero is located f(x)=-2x^4-4x^3-2x^+3x+8

OpenStudy (anonymous):

Did you try to factor the expression?

OpenStudy (anonymous):

I have no idea how to do this problem at all

OpenStudy (anonymous):

This might be a start to factor it: $$-2x^3(x+2)-2x^2+3x+8$$

OpenStudy (whpalmer4):

Can you characterize the roots you'll get?

OpenStudy (anonymous):

The possible answers are: A. there is a zero between x=1, x=2,x=-1,x=-2 B.zero between x=2, x=3,x=1,x=0,x=-2,x=-3 C.zero between x=-1,x=-2 D. zero between x=1 and x=2

OpenStudy (whpalmer4):

Oh, that takes all the challenge out of it :-)

OpenStudy (whpalmer4):

Do you know calculus?

OpenStudy (anonymous):

no im in algebra

OpenStudy (whpalmer4):

Okay, do you know about Descartes' Rule of SIgns?

OpenStudy (anonymous):

Just start to plug in those values into the equation, in the order: -2, -1, 0, 1, 2, 3 If one value is positive and the following is negative you have a zero between them.

OpenStudy (anonymous):

I don't know how to do that? I'm only in algebra one and I need this done in ten minutes to turn in :(

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

Calculate $$-2x^4-4x^3-2x^2+3x+8$$ When x= -2, then x=-1 and the other values. Example: if x=0 the answer is $$-2\times0^4 - -4\times0^3 - 0 + 8= 8$$.

OpenStudy (anonymous):

so would it be a, b, c or d

OpenStudy (anonymous):

Its not B.

OpenStudy (anonymous):

If x=1 the expression turns +3

OpenStudy (anonymous):

There is definetely a zero between 1 and 2. So its either A or D.

OpenStudy (whpalmer4):

With the equation written in descending order of powers of x, count the sign changes. That number, possibly minus a multiple of 2, is the number of positive roots. By my calculation, the answer is 1 positive root. Next, rewrite the equation substituting (-x) wherever you see x, and simplifying. For this equation, that gives -2x^4+4x^3-2x^2-3x+8. Count the sign changes again. This number, possibly reduced by a multiple of 2, gives the number of negative real roots. Here I make it to be 3 or 1. You'll always have as many total roots as the highest power exponent in the polynomial, so here there will be 4. Any not accounted for above are complex, and come in conjugate pairs. I think your best bet here is to do as @DanielM_113 suggests. I'll do a few to help: \[-2(-2)^4-4(-2)^3-2(-2)^2+3(-2)+8=-32-32-8-6+8 = -70\]\[-2(-1)^4-4(-1)^3-2(-1)^2+3(-1)+8=-2+4-2-3+8=5\]So there must be a zero between x = -2 and x = -1 You know that the value is positive at x = 0, so no root between x = -1 and x = 0. On to x=1: \[-2(1)^4-4(1)^3-2(1)^2+3(1)+8 = -2-4-2+3+8=3\]still positive, so no root between x=0 and x=1 \[-2(2)^4-4(2)^3-2(2)^2+3(2)+8 = -32-32-8+5+8 = -57\]so we have a zero between x=1 and x=2 \[-2(3)^4-4(3)^3-2(3)^2+3(3)+8 = -162-108-36+9+8 < 0\]so no zero between x=2 and x = 3

OpenStudy (anonymous):

But that wouldn't fit any of the possible answers?

OpenStudy (whpalmer4):

If you've chosen an answer, I can show you a graph of the function in the relevant area...

OpenStudy (anonymous):

And so the answer must be A.

OpenStudy (whpalmer4):

we had a zero between x=-2, x=-1, and another between x = 1, x=2

OpenStudy (anonymous):

I didnt think of doing a graph online, that is deviously genius.

OpenStudy (whpalmer4):

As I was saying about Descartes' Rule of Signs (which it sounds like you probably haven't had yet), we had 1 positive root, and either 1 or 3 negative roots. Turns out to be 1 negative root. The other two are complex numbers. The full set of solutions is: \[\{\{x\to -0.709177-1.25956 i\},\{x\to -0.709177+1.25956 i\},\]\[\{x\to -1.70468\},\{x\to 1.12303\}\}\]

OpenStudy (whpalmer4):

Though the symbolic version is quite striking :-)

OpenStudy (anonymous):

Thanks for explaining the Decartes Rule of Signs, its been some time I saw that and didnt remember it.

OpenStudy (whpalmer4):

I didn't know about it either until someone asked a question about it on OpenStudy. I had to learn quickly in order to be able to answer :-)